OpenStudy (camthemathhunter):

A telephone company charges a fixed monthly rate plus a rate per minute of usage. The company charges $120 for 100 minutes of usage and $95 for 50 minutes of usage. An equation can be written to show the relationship between the total minutes used (x) and the total monthly charges (y). Which of the following best describes the steps to draw the graph of y against x? Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 0.50 Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 2 Draw a graph which joins the points (120, 100) and (95

OpenStudy (camthemathhunter):

and has a slope = 2 Draw a graph which joins the points (120, 100) and (95, 50) and has a slope = 0.50

OpenStudy (ineedhelplz):

m= y2-y1/x2-x1 m=95-120/50-100 m=-25/-50 m=1/2

OpenStudy (ineedhelplz):

Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 0.50